OpenStudy (anonymous):

Solve by simulating the problem. You have a 5-question multiple-choice test. Each question has four choices. You don’t know any of the answers. What is the experimental probability that you will guess exactly three out of five questions correctly? Type your answer below using complete sentences.

OpenStudy (anonymous):

What exactly is giving you problems?

OpenStudy (anonymous):

the whole thing

OpenStudy (anonymous):

Well, what would the chance be to get a single question right, if we ignore the other 4?

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

The question has 4 choices, not 5.

OpenStudy (anonymous):

sorry 1/4

OpenStudy (anonymous):

Right, so what would the chance be to get 3 questions right if we ignore the other 2?

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

Well, that seems a bit strange, doesn't it? Because that would mean answering 3 questions is easier than answering one question. surely that can't be right?

OpenStudy (anonymous):

What you did here is 1/4 + 1/4 + 1/4. Try multiplying instead.

OpenStudy (anonymous):

I did 1/4*3 and I got 3/4

OpenStudy (anonymous):

True, but what I meant is \[\left( \frac{ 1 }{ 4 } \right)^3 = \frac{ 1 }{ 4}*\frac{ 1 }{ 4 }*\frac{ 1 }{4 }\] which is definitely not 3/4 Otherwise, the chance of answering 5 questions correctly would be 5/4, which would be impossible.

OpenStudy (anonymous):

1/64

OpenStudy (anonymous):

yes, that's right. Although that is still only part of our answer, since that would be the chance of getting 3 or more questions correctly, since we still ignored the other 2 questions. So we also need to take into account the two questions which must be incorrect.

OpenStudy (anonymous):

You know the chance of answering a question incorrectly (just one question on its own)?

OpenStudy (anonymous):

yes if there is only 1 question that you answer and you get it wrong its 0/1

OpenStudy (anonymous):

I mean the chance of getting the question wrong. If the chance of getting it right is 1/4 then what would the chance of getting it wrong be?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

If a question has 4 options, and only 1 of them is right, how many wrong answers are there to a question?

OpenStudy (anonymous):

3 would be wrong

OpenStudy (anonymous):

True

OpenStudy (anonymous):

So, if we use the same method as for the 3 right questions, if we want to know the chance of getting 2 wrong questions, we multiply 2 wrong questions with each other, so \[\left( \frac{ 3 }{ 4 } \right)^2 = \frac{ 3 }{ 4 }*\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

And whatever the result of that is, we multiply the outcome with \(\frac{ 1 }{ 64 }\) for the final answer.

OpenStudy (anonymous):

I got 9/256 for the answer

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

From 3/4*3/4 or the final answer?

OpenStudy (anonymous):

Becausr \[\frac{ 3 }{ 4 }*\frac{ 3 }{ 4 }=\frac{ 3*3 }{ 4*4 }=\frac{ 9 }{ 16 }\] and \[\frac{ 9 }{ 16 } *\frac{ 1 }{ 64 }=\frac{ 9*1 }{ 16*64 }=\frac{ 9 }{ 1024}\]

OpenStudy (anonymous):

Somehow your answer is 4 times larger.

OpenStudy (anonymous):

I did \[3/4^{2} * 1/16\]

OpenStudy (anonymous):

and I got 9/256

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

how what?

OpenStudy (anonymous):

How did you calculate \( 3/4^{2} * 1/16 \)

OpenStudy (anonymous):

(should be \( (3/4)^2 \) btw, because the 3 is also squared, but even then the outcome is not what i would expect.

OpenStudy (anonymous):

I would still get 9/256

OpenStudy (anonymous):

Can you do it step by step? or do oyu put it in a calculator or something?

OpenStudy (anonymous):

i put it in the calculator

OpenStudy (anonymous):

What exactly did you type into the calculator?

OpenStudy (anonymous):

\[(3/4_{2}) * 1/16\]

OpenStudy (anonymous):

That's weird because \[\left( 3/4^2 \right)*1/16 = 3/256\] and \[\left( \frac{ 3 }{4 } \right)^2*1/16 = 9/1024\] Neither is 9/256

OpenStudy (anonymous):

i keep getting 9/256

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

it's 1/64 not 1/16

OpenStudy (anonymous):

How did i not see that sooner

OpenStudy (anonymous):

So nothing is wrong with your calculator, it's just you used 1/16 instead of 1/64 which is exactly 4 times larger.

OpenStudy (anonymous):

so i took \[\left( 3/4 \right)^{2} * \left( 1/64 \right)\] i got 0.008789063

OpenStudy (anonymous):

Yeah, that's a decimal approximation, it's actually 0.0087890625 (9/1024) But it's close enough.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

You'd probably get the same value if you type 9/1024 in your calculator, maybe it only goes 10 digits. Mine goes about a 1000

OpenStudy (anonymous):

For most purposes 10 is enough

OpenStudy (anonymous):

so it is 9/1024?